OpenStudy (anonymous):

Assume that you wanted to draw a model to scale. If you drew the 1x10-4 angstrom (Å) nucleus so that its diameter was 1.5 mm, what diameter should you draw the electron cloud (assume you have a typical electron cloud with a diameter of 4 Å). Report your answer in meters to the correct number of sig figs. The Å symbol is called an “angstrom.” Chemists use this unit frequently because it’s handy to use for very small lengths. 1 = 1x10-10 meter. (It’s 1/10 the length of a nanometer.)

OpenStudy (jfraser):

If the nucleus is 1x10-4A in diameter and the electron cloud is 4A across, the ratio \[\frac{cloud}{nucleus}\] has to be the same if you're going to scale up the model.\[\frac{4A}{1x10^{-4}A} = \frac{??}{1.5mm}\]